### v2

# TeleMem: Building Long-Term and Multimodal Memory for Agentic AI

Abstract
Large language models (LLMs) excel at many NLP tasks but struggle to sustain long-term interactions due to limited attention over extended dialogue histories. Retrieval-augmented generation (RAG) mitigates this issue but lacks reliable mechanisms for updating or refining stored memories, leading to schema-driven hallucinations, inefficient write operations, and minimal support for multimodal reasoning. To address these challenges, we propose TeleMem, a unified long-term and multimodal memory system that maintains coherent user profiles through narrative dynamic extraction, ensuring that only dialogue-grounded information is preserved. TeleMem further introduces a structured writing pipeline that batches, retrieves, clusters, and consolidates memory entries, substantially improving storage efficiency, reducing token usage, and accelerating memory operations. Additionally, a multimodal memory module combined with ReAct-style reasoning equips the system with a closed-loop observe, think, and act process that enables accurate understanding of complex video content in long-term contexts. Experimental results show that TeleMem surpasses the state-of-the-art Mem0 baseline with $19\%$ higher accuracy, $43\%$ fewer tokens, and a $2.1\times$ speedup on the ZH-4O long-term role-play gaming benchmark.
footnotetext: * Equal contribution. Project lead. # Corresponding authors.
1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of natural language processing tasks [5, 37, 29]. However, their effectiveness in long-term interactions remains fundamentally constrained by the context window limitations inherent to Transformer architectures. Even though contemporary long-context models can process hundreds of thousands of tokens [10], simply enlarging the context window does not fully address the issue: as interaction rounds accumulate, models struggle to allocate attention to distant tokens, making it difficult to accurately retrieve user-specific information embedded deep in the dialogue history. As a result, they may overlook key details or even forget important personalized information over time [12], significantly hindering their practical utility in scenarios requiring sustained, long-term relationships with users.
To address the context-window limitations, retrieval-augmented generation (RAG) has become a widely adopted strategy for managing long dialogue histories. RAG systems typically encode past interactions into vector embeddings stored in external datastores and retrieve relevant information via semantic search [21, 44, 8]. Although effective at retrieving information beyond the model’s native context window, traditional RAG pipelines fundamentally lack mechanisms for updating or refining stored memories. Once written into the datastore, memories cannot be reliably modified or deleted, making it difficult to handle evolving user preferences, resolve contradictions, or maintain long-term consistency.
This inherent limitation has motivated the development of more advanced memory architectures capable of structured storage, organization, and dynamic updating of information [40, 36]. For example, MemoryBank [45] introduces a decay mechanism inspired by the Ebbinghaus forgetting curve to selectively preserve salient information. Mem0 [2] proposes a meaning-aware memory framework in which the LLM itself extracts atomic facts, retrieves semantically related memories, and decides whether to add, update, delete, or ignore them in a streaming, turn-by-turn process. Rsum [25] employs hierarchical summarization to distinguish short-term details from long-term abstractions, while Zep [22] represents dialogue history as a temporal knowledge graph to capture causal and temporal dependencies. Collectively, these systems extend RAG by incorporating notions of forgetting, abstraction, and structure, offering more principled approaches for managing long-term conversational memory.
However, despite recent progress, existing memory systems still face three major challenges in real-world deployment. First, maintaining a consistent user profile remains difficult. Current systems rely on dozens of predefined fields, while real conversations typically provide only sparse information. This mismatch forces the model to hallucinate unsupported details and leaves many fields empty, creating unnecessary computational overhead and increasing the likelihood of structural errors. Second, current memory systems exhibit low write efficiency. Each dialogue turn triggers a retrieval step and sends the top-k memories to an LLM for add, delete, update, or no-op decisions, resulting in frequent datastore writes and inefficient API round-trips that degrade throughput and latency. Third, multimodal reasoning capabilities remain limited. Most systems are predominantly text-centric and struggle to integrate or reason over memories derived from images, audio, and other modalities, significantly restricting their use in complex real-world settings.
To tackle these challenges, we propose TeleMem, a unified long-term and multimodal memory system. By extracting only dialogue-supported narrative units, TeleMem avoids schema-driven hallucinations and enables lightweight, precise updates, resulting in a compact and reliable long-term memory. To further address storage and retrieval inefficiencies, TeleMem employs a novel writing pipeline that batches summaries, retrieves related memories, clusters semantically similar entries, and applies an LLM-driven decision step before committing them to persistent storage. This process pre-aggregates and de-duplicates fragmented information, substantially improving throughput while reducing token usage and storage overhead. In addition, TeleMem incorporates a multimodal memory module that transforms raw video streams into event and object memories. Combined with a ReAct-style reasoning framework [38], this design enables a closed-loop observe–think–act process for precise reasoning over complex video content. In summary, our contributions are following:
  1. 1.
    We introduce TeleMem, a unified long-term and multimodal memory framework that leverages narrative dynamic extraction to maintain coherent, hallucination-free user profiles.
  2. 2.
    We design a structured writing pipeline that batches, retrieves, clusters, and consolidates memory entries, substantially improving storage efficiency, reducing token overhead, and accelerating memory operations.
  3. 3.
    We develop a multimodal memory module with ReAct-style reasoning, enabling end-to-end observe–think–act capabilities for complex video understanding.
  4. 4.
    Extensive experiments demonstrate that TeleMem outperforms the state-of-the-art Mem0 baseline by 19% in accuracy on the ZH-4O benchmark, while reducing token usage by 43% and achieving a $2.1\times$ speedup.
2 TeleMem System
TeleMem integrates text-based and multimodal memory into a unified long-term memory system. The text memory module includes profile memory, which captures stable user attributes through lightweight summarization and vector retrieval, and event memory, which models dynamic interaction details using summarization, retrieval, clustering, and consolidation (Sec. 2.1). Complementing these, the multimodal memory module processes video inputs through segmentation, captioning, and cross-modal embedding, and supports video-grounded reasoning via ReAct-style tool use and temporal querying (Sec. 2.2). Together, these components enable TeleMem to maintain consistent, structured, and query-efficient long-term memory across both linguistic and visual modalities.
Figure 1: Overview of the text memory writing pipeline in TeleMem. The system maintains three types of long-term memory: user profile memory $\mathcal{M}_{\mathrm{user}}$, bot profile memory $\mathcal{M}_{\mathrm{bot}}$, and event memory $\mathcal{M}_{e}$. 1). For each dialogue turn $D_{t}$, the model generates role-specific profile summaries $p_{t}^{\mathrm{user}}$ and $p_{t}^{\mathrm{bot}}$, which are written into their respective profile memory stores in parallel. 2). Event memory is constructed through a four-stage batch pipeline. Each turn $D_{t}$ is summarized into one or more textual summaries $s_{i}$, forming the set $\mathcal{S}$. For every summary $s_{i}$, the system retrieves the most relevant event memories from $\mathcal{M}_{e}$, yielding retrieval sets $R_{i}$. All summaries and retrieved items are merged into a unified candidate pool $\mathcal{U}$, which is globally clustered into semantic groups $C_{i}\in\mathcal{C}$. For each cluster $C_{i}$, entries are chronologically ordered and passed to an LLM that assigns an action $A\in\{\textit{add},\textit{delete},\textit{update},\textit{no-op}\}$, producing consolidated event memory updates that are written back into $\mathcal{M}_{e}$.
2.1 Text Memory Module
Overall, the text memory module centers around two complementary operations: memory writing, which transforms incoming dialogue into structured long-term representations through summarization, retrieval, clustering, and refinement (Sec. 2.1.1); and memory reading, which retrieves and integrates relevant profile and event information to support downstream reasoning (Sec. 2.1.2). Together, these operations enable TeleMem to continuously accumulate, organize, and utilize long-term textual knowledge in a coherent and query-efficient manner.
Formally, we denote the dialogue at turn $t$ as $D_{t}$. We maintain three memory stores: a user profile memory $\mathcal{M}_{\mathrm{user}}$, a bot profile memory $\mathcal{M}_{\mathrm{bot}}$ both capturing stable character attributes and relations from their respective perspectives, and an event memory $\mathcal{M}_{e}$, which contains dynamic, temporally grounded interaction summaries.
2.1.1 Memory Writing
The profile and event writing mechanisms enable TeleMem to construct a coherent, dynamically evolving memory space that faithfully reflects both stable role attributes and unfolding narrative developments.
Profile Memory
To capture stable, character-centric information, TeleMem constructs role-specific profile summaries from each dialogue turn $D_{t}$. Because a single turn contains contributions from both interlocutors, the system produces two separate summaries $p_{t}^{\mathrm{user}}$ and $p_{t}^{\mathrm{bot}}$, which are written into the corresponding profile memory stores $\mathcal{M}_{\mathrm{user}}$ and $\mathcal{M}_{\mathrm{bot}}$. Unlike schema-driven methods that rely on predefined fields or incremental slot updates, TeleMem imposes no fixed profile template. Instead, each $p_{t}^{\mathrm{user}}$ and $p_{t}^{\mathrm{bot}}$ is a standalone textual description grounded solely in information explicitly supported by the dialogue. This lightweight formulation avoids schema constraints, prevents unsupported inferences, and keeps profile memory flexible, grounded, and easy to retrieve throughout long-term interactions. Moreover, this role-specific formulation naturally generalizes to settings with additional participants, allowing profile memories to scale seamlessly to multi-party interactions.
Figure 2: Example JSON snippet of a text profile memory entry.
Event Memory
Event memory is responsible for capturing the dynamic, interaction-driven information that emerges throughout a multi-turn dialogue. Different from streaming methods  [2] that update memory turn by turn and often incur redundant retrievals and fragmented writes, our batch formulation enables joint processing of multiple dialogue turns, improving coherence, efficiency, and consolidation quality. Assuming that the system receives a batch of dialogue turns $\{D_{t}\}_{t=1}^{N}$. To transform these into long-term event memory, TeleMem applies a four-stage writing pipeline consisting of summarization, retrieval, clustering, and decision (Fig. 1).
  • Summarization (parallel across turns): Each dialogue turn $D_{t}$ is independently summarized into a set of one or more textual summaries, denoted as $S_{t}$. We denote the collection of all summaries produced from a batch of dialogue turns as
    $\mathcal{S}=\bigcup_{t=1}^{N}S_{t},$
  • Retrieval (parallel across summaries): For each summary $s_{i}\in\mathcal{S}$, the system performs vector retrieval from the memory store $\mathcal{M}_{e}$, obtaining the top-$k$ related memories:
    ${R}_{i}=\text{Top-}k\!\left(\text{sim}(s_{i},\mathcal{M}_{e})\right).$
  • Clustering (global, non-parallel): A unified candidate set is constructed as
    $\mathcal{U}=\left(\bigcup_{i=1}^{N}{R}_{i}\right)\cup\mathcal{S},$
    upon which global semantic clustering is applied to obtain clusters
    $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{m}\},\quad C_{i}\subseteq\mathcal{U},$
    where each cluster groups entries describing semantically related events or topics.
  • Decision (parallel across clusters): For each cluster $C_{i}$, the system orders all entries temporally and feeds them to an LLM, which determines for each item whether to perform an action $\mathcal{A}\in\{\emph{add},\emph{delete},\emph{update},\emph{no-op}\}$. This process resolves redundancies and inconsistencies and produces refined memory entries that are subsequently re-embedded and written back into $\mathcal{M}_{e}$.
Figure 3: Example JSON snippet of a text event memory entry.
2.1.2 Memory Reading
Given a query $q$, TeleMem retrieves relevant information from three memory stores: the user profile memory $\mathcal{M}_{\text{user}}$, the bot profile memory $\mathcal{M}_{\text{bot}}$, and the event memory $\mathcal{M}_{e}$. For each store, the embedding of $q$ is used to retrieve the top $5$ most similar entries. These retrieved items provide stable user attributes, stable bot attributes, and dynamic interaction context. They are then inserted into the LLM prompt as external memory, enabling retrieval-augmented reasoning that remains consistent with long-term roles and past events.
2.2 Multimodal Memory Module
Beyond long-term textual inputs, we extend the memory architecture to support multimodal video memory. Our design targets two key requirements of long-horizon video understanding: (1) retaining fine-grained visual grounding while (2) enabling efficient retrieval and reasoning over long sequences. To this end, we build a multimodal memory module that captures both event-level semantics and entity-level information (Sec. 2.2.1), and supports flexible, tool-driven access through an agentic reading mechanism (Sec. 2.2.2).
2.2.1 Memory Writing
We first divide the video stream into 10-second clips and uniformly sample two frames from each clip as the basis for memory extraction. The writing pipeline produces two complementary memory types: event memory and key–value object memory.
Event Memory
For each clip, a VLM generates a concise caption using the sampled frames. These clip-level event memories capture localized semantic events and remain fixed once written, forming a chronological record of the video $\mathcal{M}_{e}$.  [^1] We slightly overload the notation here for simplicity of presentation.
Figure 4: Example JSON snippet of a video event memory entry.
Key–Value Object Memory
For each clip, a VLM generates person and object entities and store them in a key–value structure. Each key specifies stable identity attributes including name, appearance, identity cues, and first occurrence, while each value captures the entity’s behavior or state in the clip. To respect the LLM’s context length constraints, we avoid consolidating all object entries at once. Instead, raw key–value records are accumulated and updated in batches by a text LLM, which incrementally merges redundant information and produces temporally coherent entity summaries. After all batches are processed, we obtain the final consolidated object memory $\mathcal{M}_{\mathrm{obj}}$.
Figure 5: Example JSON snippet of an object memory entry.
2.2.2 Memory Reading
We adopt a ReAct-style agent to read from the multimodal memory, where the agent iteratively selects among three tools to retrieve, refine, or verify information. These tools operate over the text-based memory stores ($\mathcal{M}_{e}$ and $\mathcal{M}_{\mathrm{obj}}$) or the original video clips (Alg. 1).
video.retrieval
This tool performs vector-based retrieval between a textual query and the text-based memory store ($\mathcal{M}_{e}$ and $\mathcal{M}_{\mathrm{obj}}$). It returns the clip indices, specifically the start and end timestamps, that correspond to the most relevant memory entries. This allows the agent to efficiently localize the video segments associated with the query.
video.rag
This tool also conducts vector retrieval against the text-based memory store, but instead of returning clip boundaries, it selects the top-$k$ most relevant memory entries and feeds them into the LLM prompt. The LLM then synthesizes an aggregated, query-grounded response using these retrieved memory elements.
video.qa
This tool performs vision-language question answering over a specified video clip. After the agent identifies a relevant clip, VQA is used to extract fine-grained visual details such as appearance attributes, spatial relations, or object states that may not be available in textual memory.
1
Input: Inital query $q$, $\texttt{LLM}(\cdot)$, $\mathcal{M}=\mathcal{M}_{e}\cup\mathcal{M}_{\mathrm{obj}}$
Output: Response
2
3Initialize history $H=[q]$;
4
5while not exceed max_iterations do
6   action, args = $\texttt{LLM}(H)$;
7
8 if action is video.retriever then
9
10    $q\leftarrow$ args ;
[…]
12
13 else if action is video.rag then
14    $q\leftarrow$ args ;
[…]
16
17 else if action is video.qa then
18    $q,t_{start},t_{end}\leftarrow\mathtt{args}$ ;
19      results $\leftarrow$ $\texttt{video.qa}(q,t_{start},t_{end})$;
20
21 else if action is finish then
22      break;
23
24
25 $H=H+[(\text{action},\text{args},\text{results})]$;
26
27
28return $\texttt{LLM}(H)$.
Algorithm 1 Multimodal Memory Reading.
3 Experiments
Dataset
We conduct experiments on two ultra-long dialogue datasets to assess agent memory capabilities. ZH-4O [1] serves as a Chinese role-playing benchmark comprising 28 authentic human-LLM sessions (avg. 600 turns), annotated with 1,068 multiple-choice probing questions to test memory recall. Complementarily, we utilize LoCoMo [16] as an English dataset, which consists of 10 long-context sessions paired with 1,540 question-answer pairs covering single-hop, multi-hop, open-domain, and temporal reasoning. Notably, we exclude the LoCoMo adversarial subset due to the absence of ground-truth answers.
Baselines
We compare our framework against five baselines, categorized into general paradigms and specialized memory systems. First, we consider two foundational approaches: Long context LLM utilizes the entire conversation history to establish a full-context reference, while RAG retrieves the top-$k$ semantically relevant segments to augment generation. Second, we include three advanced memory architectures: Memobase [24] maintains structured user profiles and event memories; A-Mem [35] utilizes an agentic framework to autonomously link memory notes; and Mem0 [2], a modular memory system designed for scalable deployment with explicit in-context memory operations.
Implementation
To ensure a fair and consistent evaluation, we re-implemented all baselines using Qwen3-8B [37] (configured in “no-think” mode) as the backbone LLM, paired with Qwen3-8B-embedding [37] for vector representations. Crucially, a unified prompt template is applied for response generation across all settings to strictly control for variance.
Evaluation
Our evaluation strategies are tailored to each dataset: for ZH-4O [1], we quantify memory fidelity via QA Accuracy, where the model selects the single correct option for each multiple-choice query against ground-truth labels. For LoCoMo [16], following prior work [2], we employ an LLM-as-a-Judge ($J$) approach using GPT-4o [5], which assesses the factual accuracy, relevance, and completeness of the generated responses.
3.1 Results on ZH-4O
Method Overall (%)
RAG 62.45
Mem0 70.20
MOOM 72.60
A-Mem 73.78
Memobase 76.78
Long context LLM 84.92
TeleMem 86.33
Table 1: Performance comparison on the ZH-4O benchmark. The metric represents QA Accuracy (%) across 1,068 probing questions. The best performance is highlighted in bold.
Main Results
Table 1 presents the performance of different memory paradigms on the ZH-4O benchmark, revealing clear distinctions in their ability to support long-horizon, multi-turn question answering. We make the following observations.
  • Retrieval-only methods are insufficient. RAG achieves the lowest accuracy (62.45%), indicating that flat semantic retrieval without temporal ordering or relational structure fails to support multi-turn, memory-intensive reasoning required by ZH-4O.
  • Explicit memory mechanisms consistently improve performance. All memory-augmented approaches outperform RAG, demonstrating the necessity of maintaining persistent and updatable memory states for long-horizon dialogue understanding. Among them, Memobase performs best (76.78%), likely due to its structured user profiling that aligns well with the role-playing characteristics of the benchmark.
  • Long-context modeling alone has inherent limitations. The Long Context LLM baseline achieves 84.92% accuracy by leveraging the full dialogue history, but its reliance on raw context makes it susceptible to noise, redundancy, and attention dilution as interactions grow longer.
  • Coordinated read–write memory yields the best results. Our proposed TeleMem attains the highest accuracy of 86.33%, outperforming both long-context and prior memory-based architectures. This demonstrates that selectively compressing salient information and enabling context-aware memory access is more effective than unstructured context accumulation or loosely coupled memory modules.
Write $\downarrow$     Read $\rightarrow$ 0.6B 1.7B 4B 8B 14B 32B Avg.
0.6B 52.72 64.23 73.69 75.47 74.06 72.19 68.73
1.7B 61.14 71.54 79.12 78.56 80.81 77.72 74.82
4B 63.48 74.25 81.09 83.33 84.55 83.33 78.34
8B 65.54 77.15 84.36 86.33 85.77 84.93 80.68
14B 67.04 78.28 84.55 86.33 87.55 85.49 81.54
32B 68.91 80.81 83.99 86.70 86.61 85.96 82.16
Avg. 63.14 74.38 81.13 82.79 83.17 81.60
Table 2: Memory read–write scaling law. Performance (QA Accuracy, %) when using different model sizes for the memory write LLM (rows) and the memory read LLM (columns). Results show consistent performance gains as either component scales up, with larger read–write combinations yielding the strongest performance.
Read-Write Scaling Law
Table 2 reveals a clear memory read–write scaling law by varying the model sizes of the memory write LLM and the memory read LLM. Overall performance improves monotonically as either component scales up, indicating that both memory writing and memory reading contribute substantially to downstream QA accuracy. However, the gains are not symmetric across the two dimensions. Increasing the size of the write LLM leads to consistent improvements across almost all read settings, suggesting that stronger writers produce more informative, compact, and robust memory representations that benefit readers of different capacities. Scaling the read LLM yields even more pronounced gains, particularly when paired with medium-to-large write models. This indicates that memory retrieval and reasoning capacity plays a critical role in exploiting stored information, especially when the memory content is sufficiently well-structured. Notably, performance saturates when the reader significantly outscales the writer, implying that retrieval capacity alone cannot compensate for low-quality memory writing. The best results are achieved when both the write and read LLMs are scaled jointly, highlighting the complementary and interdependent nature of memory writing and reading. These findings suggest that effective long-term memory systems should be designed with coordinated read–write capacity rather than over-optimizing either component in isolation.
3.2 Results on LoCoMo
Main Results
Table 3 compares different methods on the LoCoMo benchmark across four reasoning categories. Overall, the Long Context LLM achieves the strongest performance (70.71%), demonstrating the effectiveness of processing full interaction histories for complex memory-centric tasks. This advantage is particularly pronounced on temporal questions, where access to complete chronological context is crucial. In contrast, retrieval-based baselines perform poorly across all categories, highlighting the limitations of static semantic retrieval for long-term conversational reasoning. Memory systems exhibit diverse strengths across reasoning types. Mem0 performs well on single-hop questions (54.96%) and temporal reasoning (60.28%), suggesting that its modular memory abstraction is effective for fact recall and time-sensitive information. Memobase excels on multi-hop reasoning (66.04%), likely benefiting from its structured memory representations that better support compositional inference. A-Mem shows more balanced but moderate performance, indicating that agentic memory control alone is insufficient without strong memory structuring. TeleMem achieves competitive performance on single-hop (64.53%) and temporal reasoning (78.47%), approaching the Long Context LLM in these categories, which suggests that its memory abstraction effectively captures salient factual and temporal information. However, its weaker performance on multi-hop questions indicates remaining challenges in supporting complex relational reasoning over stored memories. Overall, these results highlight that different memory mechanisms favor different reasoning patterns, and no single approach uniformly dominates all categories, underscoring the importance of task-aware memory design for long-horizon QA.
Method Single-Hop Multi-Hop Open Domain Temporal Overall
RAG 20.91 32.39 35.41 48.03 39.03
Mem0 54.96 31.15 40.62 60.28 52.01
A-Mem 44.32 35.82 33.33 56.59 48.57
Memobase 50.70 66.04 37.50 58.97 57.59
Long context LLM 64.89 47.66 42.70 84.66 70.71
TeleMem 64.53 20.56 40.62 78.47 61.49
Table 3: Performance comparison on the LoCoMo benchmark. All metrics represent accuracy (%) evaluated via LLM-as-a-Judge across 1,540 question-answer pairs.
4 Related Work
Text LTM
Recent research on equipping Large Language Models (LLMs) with text long-term memory has primarily evolved through system-level abstractions, cognitive-inspired mechanisms, and structured agentic frameworks. To transcend fixed context windows, operating system paradigms like MemGPT [20] and MemOS [7] leverage virtual memory abstractions and hierarchical scheduling to orchestrate data flow between active context and external storage. Complementing these architectural innovations, bio-inspired models such as MemoryBank [45], LightMem [3], and Nemori [19] integrate human cognitive theories—ranging from the Ebbinghaus forgetting curve to the Atkinson-Shiffrin model—to dynamically prioritize, segment, and decay information for efficient retrieval. On the structural front, Mem0 [2], A-Mem [35], and MIRIX [30] propose modular systems that enable self-evolving knowledge consolidation and multi-agent coordination.
Graph LTM
Recent literature on long-term memory for LLM agents highlights a transition from unstructured storage to graph-based architectures that enhance reasoning, consistency, and scalability. Foundational works like ${Mem0}^{\tiny g}$ [2] and Zep [22] establish production-ready and temporal knowledge graph systems to manage agent memory efficiently. To address the complexity of long-term interactions, G-Memory [41], LiCoMemory [4], and SGMEM [32] proposes hierarchical graph structures that organize information at varying granularities, from sentence-level details to high-level insights, thereby facilitating self-evolution and reducing fragmentation. Enhancing the reasoning capabilities over these structures, GraphCogent [26] and D-SMART [6] integrate working memory models and dynamic reasoning trees to support complex graph understanding and dialogue consistency, MemQ [34] focuses on optimizing knowledge graph reasoning. Furthermore, [33] introduces trainable graph memories that abstract agent trajectories into strategic meta-cognition. Collectively, these studies demonstrate that structured, hierarchical, and dynamic graph memories are essential for developing autonomous agents capable of coherent, long-term strategic planning and reasoning.
Parametric LTM
Parametric memory aims to encode knowledge or contextual information directly into the weights or persistent hidden states of neural networks [14, 13, 18], distinguishing itself from non-parametric approaches that rely on external vector databases. Wang et al. [27] propose a decoupled memory mechanism, employing a residual side network to cache long-term context while keeping the base LLM frozen. Similarly, MLPMemory [31] internalizes retrieval by training multilayer perceptrons to approximate k-nearest neighbor distributions as differentiable mappings. Pushing the internalization of retrieval further, Tay et al. [23] introduce the Differentiable Search Index, which eliminates external indices by training the model to map queries directly to document identifiers via its parameters. regarding personalized adaptation, Zhang et al. [42] combine LoRA with Bayesian optimization to inject dialogue history into model weights. Complementary to these injection methods, Meng et al. [17] explore the interpretability of parametric storage, proposing ROME to locate and directly edit factual associations within the Transformer’s MLP layers. Overall, parametric memory methods offer scalability and unified reasoning but face challenges regarding update costs and potential misalignment with the base model.
Multimodal LTM
Multimodal memory [28, 43] has drawn increasing attention as modern agents must store and reason over long-horizon visual and textual information. Mem0 [2] introduces a multimodal image interface but ultimately reduces visual inputs to captions and continues to operate purely in the textual space. M3-Agent [15] extends memory into the multimodal domain by maintaining entity-level representations across audio–visual streams and enabling agentic retrieval over long video sequences. Inspired by hippocampal mechanisms, HippoMM [9] proposes cross-modal event encoding and temporal consolidation to support richer multimodal recall. MemVerse [11] further explores lifelong multimodal memory through a hierarchical retrieval framework with periodic distillation for compactness. In contrast, VisMem [39] emphasizes preserving visual latent representations rather than relying solely on textual abstractions. Overall, existing multimodal memory systems highlight the need for multi-granular representations and adaptive retrieval, while our approach advances this direction with a lightweight hierarchical design tailored for continuous multimodal streams.
5 Conclusion
We introduced TeleMem, a unified long-term and multimodal memory system that overcomes key limitations of existing RAG-based approaches. By extracting narrative-grounded information and employing a structured writing pipeline for batching, clustering, and consolidation, TeleMem maintains coherent user profiles while greatly improving storage and token efficiency. Its multimodal memory module with ReAct-style reasoning further enables accurate observe–think–act processing for complex video content. Experiments on the ZH-4O benchmark show that TeleMem substantially outperforms Mem0 in accuracy, efficiency, and speed, underscoring its effectiveness.
Acknowledgments.
The development of TeleMem has been greatly influenced by the contributions of open-source communities and prior research efforts. We express our sincere appreciation to the teams and projects whose work has inspired and supported this system, including Mem0, Memobase, MOOM, DVD, and Memento. Their innovations have provided valuable foundations upon which this work builds.